OpenStudy (anonymous):

If J = 40° and G = 60°, then HJ _____ GJ Choose the correct symbol that would make a true statement. < = >

OpenStudy (anonymous):

well what is H? unless we know that then we cant honestly tell you

OpenStudy (anonymous):

yu wanna go on the wesite and do it for me Bsho

OpenStudy (anonymous):

no i cant do it for you

OpenStudy (anonymous):

you have to do your own work but can try to help

OpenStudy (anonymous):

idk wat H is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait is it about a triangle?

OpenStudy (anonymous):

cuz if thats so then H would be 80

OpenStudy (anonymous):

yea its about a triangle

OpenStudy (anonymous):

ok then let me look.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im having trouble so if anyone beats me to it then good. its been 2 years since i was in geomatry and we didnt do alot on this topic

OpenStudy (anonymous):

well if the angles are added then HJ would be greater then GJ becasue 80 + 40 =120 and 40+60=100. so im going to go with greater then. your third choice

OpenStudy (anonymous):

okay, j, g, and h all add together to make 180. you know that j = 40 and g = 60, so add both of those together and you get 100. so from common sense, you know h would be 80. here's what your problem is asking: \(is~H+J~more,~less,~or~equal~to~G+J?\) well, plug in your answers. \(80+40=120\) \(60+20+100\) \(120~O~100\) is 120 more or less that 100? (hint: <)

OpenStudy (anonymous):

nononon, >

OpenStudy (anonymous):

stupid me :P

OpenStudy (anonymous):

thanx elven. i was being stupid and trying to mulitply them. then i rememeber that angles are added not mutiplied. but theres your answer @mikerip

OpenStudy (anonymous):

thx ;P

OpenStudy (anonymous):

*:P

OpenStudy (anonymous):

I know i did geomatry 2 years ago and trig this year but my trig class was mixed in with pre calc so we didnt go over alot of recap

OpenStudy (anonymous):

haha, what grade you in?

OpenStudy (anonymous):

im a junior in highschool

OpenStudy (anonymous):

i'm in 7th